(a^2+b^2)\(ab+1) find an equation for solving
1
Expert's answer
2019-01-11T12:31:40-0500

An equation (a2+b2)/(ab+1)=0 has a solution if and only if the numerator is 0, and the denominator is not 0.

The numerator is 0 if and only if a = b = 0 because the sum of the squares of two numbers is 0 if and only if each number is 0.

In this case, the denominator ab +1=0+1=1 is not 0.

Answer: a=0, b=0.